OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0 7-3x^2=(x+2)^2-6

OpenStudy (amistre64):

expand and combine

OpenStudy (amistre64):

(x+2)^2 expands into what?

OpenStudy (anonymous):

x^2+4

OpenStudy (amistre64):

very close; good effort; but lets do this the long way to see what we get... (x+2) (x+2) ; distribute it out once like this: (x+2)(x) + (x+2)(2) ; now distribute it thru again like this: xx + 2x + 2x + 4 ; now combine the terms that look the same x^2 +4x +4

OpenStudy (amistre64):

now lets put this exapnded form to good use: 7 -3x^2 = (x^2 +4x +4) -6 7 -3x^2 = x^2 +4x +4 -6 put it all to one side; i choose right. 0 = -7 +3x^2 +x^2 +4x +4 -6 ; add everything that looks the same 0 = 4x^2 + 4x -9 right?

OpenStudy (anonymous):

yep thats right

OpenStudy (amistre64):

yay!! ... i do mess up alot and its nice to have confirmation :)

OpenStudy (anonymous):

no problem